# Dynamics of a Multicomponent Dark Energy Model and the Possibility of EDE like behaviour

### v1

@@ Abstract
This study explores the dynamics and phase-space behavior of a multicomponent dark energy model, where the dark sector consists of a minimally coupled canonical scalar field and the cosmological constant, using a dynamical system analysis setup for various types of potentials for which a general parameterisation of the scalar field potentials has been considered. Several fixed points with different cosmological behaviours have been identified. A detailed stability analysis has been done and possible late-time attractors have been found. For this multi-component dark energy model, the late-time attractors are either fully dominated by the cosmological constant or represent a scenario where a combination of the scalar field and the cosmological constant dominates the universe. We have also shown that for this type of model, the scalar field can show early dark energy (EDE) like behaviour. However, our analysis indicates that this EDE like behaviour occurs naturally deep in the matter-dominated era, not before the recombination era.
@@ I Introduction
Various dynamical dark-energy models have been suggested as alternatives for the cosmological constant [17, 18]. For dynamical dark energy models the equation of state of the dark energy changes over time [19, 20, 21, 22, 23, 24, 25, 26, 27]. These models include but are not limited to quintessence, k-essence, and phantom-type scalar field models, where generally a scalar field is coupled with the matter minimally or non-minimally with a associated potential which can generate sufficient negative pressure to drive the accelerated expansion of the universe. Recent observations from DESI collaboration [28, 29] and other literatures [30, 31, 32, 33] have pointed towards the evidence for the dynamical nature of the dark energy over the cosmological constant.
@@
Recent literature also shows evidence towards the possibility of a multi-component nature of the dark energy sector. These models are more preferable compared to the cosmological constant when aligned with current observations [34, 27, 35, 36, 37]. In these types of models, the dark sector is considered to consist of cosmological constant together with other dark energy components like a scalar field[34] or a fluid[35], or to be composed of multi-scalar fields[27]. In addition, these models can significantly reduce Hubble tension compared to the $\Lambda$CDM model[34, 27].
@@
In this paper, we have investigated a multicomponent dark energy model where the dark sector comprises a minimally coupled canonical scalar field, known as a quintessence field, and a cosmological constant. Our primary goal is to examine the phase space behaviour and dynamics of this composite model using the dynamical systems approach. By applying appropriate variable transformations, we convert the system of equations into a set of autonomous equations. Subsequently, these equations are recast in polar form to facilitate mathematical handling. We have considered a parameterization of the scalar field potential that can encompass various forms of potential to keep our analysis as general as possible, as there is no consensus on the choice for the form of the scalar field potential. A detailed stability analysis of the system is presented. Additionally, we explore the dynamics of this model by numerically solving it, revealing some naturally occurring Early Dark Energy (EDE) like behaviours of the scalar field.
@@
The manuscript is structured as follows. In Section II, we cover the mathematical setup of the model. Section III forms the autonomous system, which is then transformed into polar form in Section IV. Section V provides a detailed stability analysis of the system. In Section VI, we explore the EDE like behaviour of the scalar field with various potentials. Finally, in Section VII, we summarise and conclude our findings.
@@ III The Dynamical System
Thus, for a particular choice of potential, there will be a corresponding form of $f$. It is also possible to consider a particular form of the $f$ and find the corresponding potential using the definition of the $y,\lambda$ and $\Gamma$ variable given in Eq.(6). In this work, we consider $f=\alpha_{0}+\alpha_{1}\lambda+\alpha_{2}\lambda^{2}$, where the parameters $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ can take any real value. This particular parametrisation of $f$ was first proposed in [38]. Different choices of the $\alpha_{0},\alpha_{1}$ and $\alpha_{2}$ parameters will correspond to different classes of potentials. In Table-1, we list eight distinct classes of potentials, each corresponding to different choices of the $\alpha$ parameters. The classification was made based on whether certain $\alpha$ parameters are zero or not. This approach has been taken to facilitate our analysis for a wide class of potentials in a single setup. A similar approach but for a different choice of dynamical variables has also been used in [39].
@@ V.1 Radiation Dominated Era
The equilibrium points indicating the radiation domination era are characterised by $\Omega_{rc}=1$ and represented by ‘$r$’ which are of the nonisolated type and given in Table-2. For all the equilibrium points of this era, the corresponding Jacobian matrix has two different sets of eigenvalues depending on whether $\theta_{c}=n\pi$ or $\theta_{c}=n\pi\pm\frac{\pi}{2}$. Those eigenvalues are listed in Table-3. In both cases, the Jacobian matrix has both positive and negative eigenvalues. So, these equilibrium points are saddle-like in nature. These equilibrium points exist for all potentials ($V_{1}-V_{8}$) listed in Table-1. From Table-4 it can be seen as expected that at these equilibrium points the universe is decelerating, and $w_{tot}=1/3$.
@@ V.2 Matter Dominated Era
The equilibrium points indicating the matter dominated era are defined by $\Omega_{mc}=1$, are denoted by ‘$m$’ and given in the Table-2. The Jacobian matrix, corresponding to any equilibrium point during this era, exhibits two distinct sets of eigenvalues depending on whether $\theta_{c}=n\pi$ or $\theta_{c}=n\pi\pm\frac{\pi}{2}$. These eigenvalues are detailed in Table-3. For each $n\in\mathbb{Z}$, the equilibrium points during this period are categorized as saddle points due to the presence of both positive and negative eigenvalues corresponding to these fixed points. The equilibrium points for the matter domination era are present for all potentials $V_{1}-V_{8}$. One can see from Table-4 that these points represent a decelerating universe with $w_{tot}=0$.
@@ V.3 Scalar field dominated era
For the equilibrium points in the scalar field domination era, we have $r_{c}=1$. In this case, the equilibrium points are given by $e1$ $e4_{\pm}$ in Table-2. The eigenvalues of the Jacobian matrix at $e1$ depend on $\alpha_{1}$. But, for the existence of positive eigenvalues of the Jacobian matrix at $e1$, the equilibrium points given by $e1$ are unstable and exist for the potentials $V_{2}$ and $V_{6}-V_{8}$. These equilibrium points represent a decelerating universe with $w_{tot}=1$.
The equilibrium points indicated by $e2$ are of nonhyperbolic nature, characterised by a zero eigenvalue. These fixed points may be stable within the parameter range $0<\alpha_{0}\leq\frac{3}{4}$, as the remaining two eigenvalues can be negative. Due to their non-hyperbolic nature, linear stability analysis cannot determine their stability. Instead, stability must be assessed using centre manifold theory or by numerically plotting the system’s phase around the fixed point. Given that the system’s dimension exceeds three, a complete phase plot is impractical; hence, phase portraits are drawn on various projected planes. If all projected phase planes depict the equilibrium point as stable, it is stable in the full space. Conversely, if any projected phase space shows instability, the equilibrium point is unstable in the entire space. For the equilibrium point $e2$, we draw a phase plot (Fig.1) in the ($r,\Omega_{m}$) plane. This phase plot indicates that the equilibrium points specified by $e2$ are unstable on the ($r,\Omega_{m}$) plane and consequently in the full space. Those equilibrium points represent an accelerating universe with $w_{\phi}=w_{tot}=-1$.
@@
The equilibrium points, labelled $e3$ and $e4_{\pm}$, are identified for $r_{c}=1$, $\Omega_{mc}=0$, $\Omega_{rc}=0$, and $\lambda_{c}=A_{\pm}$ (refer to Appendix: A for the complete form of $A_{\pm}$). These fixed points exist for the potentials $V_{1}-V_{3}$ and $V_{7}$. However, the values of $\theta_{c}$ are not the same for these points. These equilibrium points exist when $\alpha_{2}\neq 0$ and $\alpha_{1}^{2}-4\alpha_{0}\alpha_{2}\geq 0$. The eigenvalues of $e3$ are determined by whether $n$ is even or odd, while the eigenvalues of $e4_{\pm}$ are independent of $n$. In all cases, the Jacobian matrix evaluated at $e3$ and $e4_{\pm}$ exhibits positive eigenvalues. Consequently, the fixed points represented by $e3$ and $e4_{\pm}$ are unstable regardless of the values of $n$ and the parameters $\alpha_{0},\alpha_{1}$, and $\alpha_{2}$. These fixed points represent a decelerating universe with $w_{\phi}=w_{tot}=1$.
@@ V.4 $\Lambda$ Dominated Era
The equilibrium points representing the cosmological constant or $\Lambda$-domination era are represented by $r_{c}=0$, $\Omega_{mc}=0$, and $\Omega_{rc}=0$. These fixed points exist for all potentials listed in Table-1. In this case, the equilibrium points are given by ‘$c$’ in the Table-2 and they are nonisolated equilibrium points. The eigenvalues of the Jacobian matrix evaluated at ‘$c$’ depend on $\theta_{c}$. For $\theta_{c}=n\pi$ the Jacobian matrix has a mixture of positive and negative eigenvalues, hence the fixed points in this case are saddle. On the other hand, finding the stability of the equilibrium points ‘$c$’ for $\theta_{c}=n\pi\pm\frac{\pi}{2}$ is more involved since it has two zero eigenvalues and other all negative eigenvalues. Three negative eigenvalues correspond to a 3D stable manifold, whereas the two zero eigenvalues correspond to a 2D center manifold. In general, one can use the center manifold theorem to investigate the stability of this fixed point, rather here we draw the phase plots on 3D planes. All the points in red color and the points on the lines shown in red in the projected phase spaces, as depicted in Fig. 2, represent the equilibrium point ‘$c$’ for the $V_{2}$ potential. From this figure, we can see that the fixed points given by ‘$c$’ are stable for $\theta=n\pi\pm\frac{\pi}{2}$ in all projected subspaces and hence stable in the entire 5D space. We have also checked that for the other class of potentials, the behaviour of the fixed point remains the same. These fixed points represent an accelerating universe with $w_{\phi}=1$ and $w_{tot}=-1$ for $\theta=n\pi$ whereas $w_{\phi}=w_{tot}=-1$ for $\theta=n\pi\pm\frac{\pi}{2}$. Interestingly, in the first case, the scalar field behaves as a stiff fluid.
Figure 2: Phase plots in different $3D$ projected spaces for the equilibrium point ‘$c$’ and for the potential $V_{2}$.
Figure 3: Plots of the numerical solutions for each dynamical variable by considering perturbations around the equilibrium point $ec1$. The plots in the left column are shown for potential $V_{1}$ and on the right it is shown for the potential $V_{8}$.
@@ V.5 The Era Shared by Scalar Field and other components
The fixed points $er1_{\pm}$ to $er6_{\pm}$ represent the state where the universe’s energy budget is shared by radiation and the scalar field by the relation $r_{c}^{2}+\Omega_{rc}=1$. These equilibrium points are detailed in the Table-2. For each $n\in\mathbb{Z}$, the equilibrium points $er1_{\pm}$, $er2_{\pm}$, $er5_{\pm}$, and $er6_{\pm}$ exhibit positive eigenvalues, making these points unstable or saddle depending on the choice of parameters $\alpha_{0},\alpha_{1},\alpha_{2}$. The eigenvalues of the Jacobian matrix at $er3_{\pm}$ and $er4_{\pm}$ are complicated in general, making it difficult to draw any conclusions about their stability. For these points, we evaluate the stability at specific values of $\alpha$ parameters. We have considered the same value of the $\alpha$ parameters that have been used to plot Fig.4 and Fig.6, for potentials $V_{1}$ and $V_{3}$ respectively. For the potential $V_{1}$ the eigenvalues at $er3_{\pm}$ are $1,4,-2.8394$, and $-0.5\pm 1.93041i$, while at $er4_{\pm}$ they are $-201.793,-11.2262,10.2262,4$, and $1$, indicating that these points are saddle points. For potential $V_{3}$, the eigenvalues at $er3_{\pm}$ and $er4_{\pm}$ are $1,4,-4.48$, and $-0.5\pm 1.92544i$, showing that they are saddle points. The fixed points $er1_{\pm}$ and $er2_{\pm}$ exist for the potentials $V_{1}$ and $V_{3}$; $er5_{\pm}$ exist for the potential $V_{2}$; and $er6_{\pm}$ for potential $V_{4}$. For all possible equilibrium points in this era $w_{\phi}=w_{tot}=1/3$ representing a decelerated universe.
@@
For the equilibrium points in the era shared by scalar field and matter, we have $r_{c}^{2}+\Omega_{mc}=1$. In this case, the equilibrium points are given by $em1_{\pm}$ $em7$ and are given in Table-2. For each $n\in\mathbb{Z}$, though the eigenvalues of the Jacobian matrix evaluated at the equilibrium points $em1_{\pm}$, $em2_{\pm}$ and $em4-em7$ depend on the parameters $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$; but it has a positive eigenvalue irrespective of $\alpha_{i}$. So, the equilibrium points given in this era are unstable in nature. The equilibrium points $em1_{\pm}-em3_{\pm}$ exist for the potentials $V_{1}$, $V_{3}$; $em4$ and $em5$ exist for the potential $V_{2}$; $em6$ and $em7$ exist for the potential $V_{4}$. The eigenvalues of the Jacobian matrix evaluated at the equilibrium points $em3_{\pm}$ are very complicated. So, for $em3_{\pm}$, we evaluate the stability similarly as the case for $er3_{\pm}$ and $er4_{\pm}$. For the potential $V_{1}$ the eigenvalues at $em3_{+}$ are $3,-1,-2.12955$, and $-0.75\pm 1.97931i$, while at $er3_{-}$ they are $9.06334,3,-1,-10.5633$, and $-151.345$, indicating that these points are saddle points. For potential $V_{3}$, the eigenvalues at $em3_{\pm}$ are $3,-1,-3.36$, and $-0.75\pm 1.97522i$, showing that they are saddle points. The exact value of the deceleration parameter depends on $\alpha$ parameters for all the potentials belonging to this era. However, for all possible equilibrium points in this era $w_{\phi}=w_{tot}=0$ representing a decelerated universe.
@@
For the equilibrium points in the era shared by scalar field and $\Lambda$, we have $r_{c}^{2}+\Omega_{\Lambda_{c}}=\Omega_{\phi_{c}}+\Omega_{\Lambda_{c}}=1$. In this case, the equilibrium points are given by $ec1$ and $ec2_{\pm}$. The equilibrium points given by $ec1$ exist for all potentials $V_{1}-V_{8}$. Those equilibrium points are given in the Table-2. The nature of the equilibrium points given by $ec1$ depends on the parameter $\alpha_{0}$. If $0<\alpha_{0}<\frac{3}{4r_{c}^{2}}$, then the equilibrium points given by $ec1$ are normally hyperbolic equilibrium points and the Jacobian matrix has four eigenvalues with negative real part and one zero eigenvalues. Again, if $\alpha_{0}=0$ then the equilibrium points given by $ec1$ are of non-hyperbolic type and the Jacobian matrix has three negative and two zero eigenvalues. To check the stability of those equilibrium points we plot the evolution of the dynamical variables $r,\theta,\Omega_{m},\Omega_{r}$ and $\lambda$ by solving the system numerically where the system has been perturbed from the equilibrium point $ec1$. This approach for finding the stability of the non-hyperbolic fixed points has been previously used in [41, 42, 43]. In Fig.3 for example we have plotted the evolution of the system with $N$ for the potentials $V_{1}$ (on the left column) and $V_{8}$ (on the right column) subjected to perturbations around the fixed point $ec_{1}$. We can see that the evolution of the dynamical variables does not diverge but rather stays in the neighbourhood of the fixed point, indicating its stable nature. We have also checked that the qualitative behaviour of the fixed point remains the same even for the other potentials. Thus, the fixed points given by $ec1$ are stable for $\alpha_{0}=0$. Again, if $\alpha_{0}<0$, then the equilibrium points given by $ec1$ are non-hyperbolic type and the Jacobian matrix has a positive eigenvalue. Thus, the equilibrium points given by $ec1$ are stable for $0\leq\alpha_{0}\leq\frac{3}{4r_{c}^{2}}$, where $0\leq r_{c}\leq 1$ and unstable for $\alpha_{0}<0$. These fixed points represent an accelerating universe with $w_{\phi}=w_{tot}=-1$.

### v3

# Quintessence scalar field and cosmological constant: Dynamics of a multi-component dark energy model

@@ Abstract
This study explores the dynamics and phase-space behavior of a multi-component dark energy model, where the dark sector consists of a minimally coupled canonical scalar field and the cosmological constant, using a dynamical system analysis setup for various types of potential for which a general parameterization of the scalar field potentials has been considered. Several fixed points with different cosmological behaviors have been identified. A detailed stability analysis has been done and possible late-time attractors have been found. For this multi-component dark energy model, the late-time attractors are either fully dominated by the cosmological constant or represent a scenario where a combination of the scalar field and the cosmological constant dominates the universe. In this type of model, there is a possibility that the scalar field can become dynamical quite early compared to the standard era of dark energy domination. However, our analysis indicates that this early time contribution of the scalar field occurs deep in the matter-dominated era, not before the recombination era.
@@ I Introduction
Various dynamical dark-energy models have been suggested as alternatives for the cosmological constant [17, 18]. In dynamical dark energy models the equation of state of the dark energy changes over time [19, 20, 21, 22, 23, 24, 25, 26, 27]. These models include but are not limited to quintessence, k-essence, and phantom-type scalar field models, where generally a scalar field is coupled with the matter minimally or non-minimally with a associated potential which can generate sufficient negative pressure to drive the accelerated expansion of the universe. Recent observations from DESI collaboration [28, 29] and other literature [30, 31, 32, 33] have pointed towards the evidence for the dynamical nature of the dark energy over the cosmological constant.
@@
The unknown nature of dark energy leads to the question of whether the dark sector consists of a single or multiple components. Recent literature [34, 27, 35, 36, 37] also shows that the multi-component nature of the dark energy sector might be more preferable by current observations than the cosmological constant. Current observations also suggest that the equation of state of the scalar field might have a phantom barrier $(w_{\phi}=-1)$ crossing in the recent past [28, 31, 30]. In [38, 36] it has been argued that there exists a no-go theorem which prohibits a dark energy model with a single degree of freedom to cross the phantom barrier. A multi-component model of dark energy is needed for the successful phantom barrier crossing. In these types of multi-component models, the dark sector is considered to consist of a cosmological constant together with other dark energy components like a scalar field [34] or a fluid [35], or to be composed of multi-scalar fields [27]. In [34] a multi-component dark energy model has been proposed with a phantom scalar field and a cosmological constant, in [35] cosmological constant is considered with a fluid as dark energy and in [27] a generic form of the quintom model has been considered with quintessence and a phantom field as the component of dark energy. In addition, all these works suggest that multi-component models fit the data better than single component models and can significantly reduce Hubble tension compared to the $\Lambda$CDM model.
@@
One can study the phenomenology of these multi-component dark energy models by comparing them against the state of the art cosmological observations. One can also use the powerful techniques of dynamical system analysis to investigate the phase-space behavior of these models without finding an exact solution. Dynamical system analysis is particularly useful when we are interested in the asymptotic behaviour of the system. Since these models have multiple components, it is crucial to understand their late-time behavior to identify which component might dominate the universe as the late-time attractor. In this paper, we use dynamical system analysis to investigate a multi-component dark energy model where the dark sector comprises a minimally coupled canonical scalar field, known as a quintessence field, and a cosmological constant. Our primary goal is to examine this composite model’s phase space behaviour and dynamics using the dynamical systems approach. By applying appropriate variable transformations, we convert the system of equations into a set of autonomous equations. Subsequently, these equations are recast in polar form to facilitate mathematical handling. We have considered a parameterization of the scalar field potential that can incorporate various forms of potential to keep our analysis as general as possible since there is no consensus on the choice for the form of the scalar field potential. A detailed stability analysis of the system is presented. Additionally, we explore the dynamics of this model by numerically solving it, revealing the possibility of early time contribution of the scalar field.
@@
The manuscript is structured as follows. In Section II, we cover the mathematical setup of the model. Section III forms the autonomous system, which is then transformed into polar form in Section IV. Section V provides a detailed stability analysis of the system. In Section VI, we numerically investigated the evolution of the model for different classes of potentials. Finally, in Section VII, we summarise and conclude our findings.
@@ III The Dynamical System
In one approach, a specific potential is selected, determining the corresponding form of $f$. This method has already been utilized in the literature; for examples, see [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. Conversely, one can start with a specific form of $f$ and derive the corresponding potential using the definitions of $y$, $\lambda$, and $\Gamma$ variables as given in Eq.(6). To ensure our analysis applies to a wide class of potentials, we adopt the form $f=\alpha_{0}+\alpha_{1}\lambda+\alpha_{2}\lambda^{2}$, where $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$ are real-valued parameters. This particular parametrisation of $f$ was first proposed in [74]. Different choices of the $\alpha_{0},\alpha_{1}$ and $\alpha_{2}$ parameters will correspond to different classes of potentials. In Table-1, we list eight distinct classes of potentials, each corresponding to different choices of the $\alpha$ parameters. The classification was made based on whether certain $\alpha$ parameters are zero or not. This approach has been taken to facilitate our analysis for a wide class of potentials in a single setup. A similar approach, but for a different choice of dynamical variables, has also been used in [75]. Within the potentials listed in Table-1, $V_{3}$ and $V_{8}$ has been already studied in [76] and [77] respectively for early time solutions. Whereas in [78] dynamical system analysis for $V_{7}$ has been studied for the quintessence field.
@@ V.1 Radiation Dominated Era
The equilibrium points indicating the radiation domination era are characterised by $\Omega_{rc}=1$ and represented by ‘$r$’ which are of the nonisolated type. For all the equilibrium points of this era, the corresponding Jacobian matrix has two different sets of eigenvalues depending on whether $\theta_{c}=n\pi$ or $\theta_{c}=n\pi\pm\frac{\pi}{2}$. In both cases, the Jacobian matrix has both positive and negative eigenvalues. So, these equilibrium points are saddle-like in nature and exist for all potentials ($V_{1}-V_{8}$) listed in Table-1. From Table-4 it can be seen as expected that at these equilibrium points the universe is decelerating, and $w_{tot}=1/3$.
@@ V.2 Matter Dominated Era
The equilibrium points indicating the matter dominated era are defined by $\Omega_{mc}=1$, and are denoted by ‘$m$’. The Jacobian matrix, corresponding to any equilibrium point during this era, exhibits two distinct sets of eigenvalues depending on whether $\theta_{c}=n\pi$ or $\theta_{c}=n\pi\pm\frac{\pi}{2}$. For each $n\in\mathbb{Z}$, the equilibrium points during this period are categorized as saddle points due to the presence of both positive and negative eigenvalues corresponding to these fixed points. These equilibrium points for the matter domination era are present for all potentials $V_{1}-V_{8}$ and represent a decelerating universe with $w_{tot}=0$.
@@ V.3 Scalar field dominated era
For the equilibrium points in the scalar field domination era, we have $r_{c}=1$. In this case, the equilibrium points are given by ‘$e1$’ - ‘$e4_{\pm}$’. The eigenvalues of the Jacobian matrix at ‘$e1$’ depend on $\alpha_{1}$. But, for the existence of positive eigenvalues of the Jacobian matrix at ‘$e1$’, the equilibrium points given by ‘$e1$’ are unstable and exist for the potentials $V_{2}$ and $V_{6}-V_{8}$. These equilibrium points represent a decelerating universe with $w_{tot}=1$.
The equilibrium points indicated by ‘$e2$’ are nonhyperbolic, characterised by a zero eigenvalue. These fixed points may be stable within the parameter range $0<\alpha_{0}\leq\frac{3}{4}$, as the remaining two eigenvalues can be negative. Due to their non-hyperbolic nature, linear stability analysis cannot determine their stability. Instead, stability must be assessed using center manifold theory or by numerically plotting the system’s phase around the fixed point. Given that the system’s dimension exceeds three, a complete phase plot is impractical; hence, phase portraits are drawn on various projected planes. If all projected phase planes depict the equilibrium point as stable, it is stable in the full space. Conversely, if any projected phase space shows instability, the equilibrium point is unstable in the entire space. For the equilibrium point ‘$e2$’, we draw a phase plot (Fig.1) in the ($r,\Omega_{m}$) plane. This phase plot indicates that the equilibrium points specified by ‘$e2$’ are unstable on the ($r,\Omega_{m}$) plane and consequently in the full space. Those equilibrium points represent an accelerating universe with $w_{\phi}=w_{tot}=-1$.
@@
The equilibrium points, labelled as ‘$e3$’ and ‘$e4_{\pm}$’, are identified for $r_{c}=1$, $\Omega_{mc}=0$, $\Omega_{rc}=0$, and $\lambda_{c}=A_{\pm}$ (refer to Appendix: A for the complete form of $A_{\pm}$). These fixed points exist for the potentials $V_{1}-V_{3}$ and $V_{7}$. However, the values of $\theta_{c}$ are not the same for these points. These equilibrium points exist when $\alpha_{2}\neq 0$ and $\alpha_{1}^{2}-4\alpha_{0}\alpha_{2}\geq 0$. The eigenvalues of ‘$e3$’ are determined by whether $n$ is even or odd, while the eigenvalues of ‘$e4_{\pm}$’ are independent of $n$. In all cases, the Jacobian matrix evaluated at ‘$e3$’ and ‘$e4_{\pm}$’ exhibits positive eigenvalues. Consequently, the fixed points represented by ‘$e3$’ and ‘$e4_{\pm}$’ are unstable regardless of the values of $n$ and the parameters $\alpha_{0},\alpha_{1}$, and $\alpha_{2}$. These fixed points represent a decelerating universe with $w_{\phi}=w_{tot}=1$.
@@ V.4 $\Lambda$ Dominated Era
The equilibrium points representing the cosmological constant or $\Lambda$-domination era are represented by $r_{c}=0$, $\Omega_{mc}=0$, and $\Omega_{rc}=0$ and exist for all classes of potentials. In this case, the equilibrium points are given by ‘$c$’ in the Table-2 and they are nonisolated equilibrium points. The eigenvalues of the Jacobian matrix evaluated at ‘$c$’ depend on $\theta_{c}$. For $\theta_{c}=n\pi$ the Jacobian matrix has a mixture of positive and negative eigenvalues, hence the fixed points in this case are saddle. On the other hand, finding the stability of the equilibrium points ‘$c$’ for $\theta_{c}=n\pi\pm\frac{\pi}{2}$ is more involved since it has two zero eigenvalues and other all negative eigenvalues. Three negative eigenvalues correspond to a 3D stable manifold, whereas the two zero eigenvalues correspond to a 2D center manifold. In general, one can use the center manifold theorem to investigate the stability of this fixed point, rather here we draw the phase plots on 3D planes. All the points in red color and the points on the lines shown in red in the projected phase spaces, as depicted in Fig. 2, represent the equilibrium point ‘$c$’ for the $V_{2}$ potential. From this figure, we can see that the fixed points given by ‘$c$’ are stable for $\theta=n\pi\pm\frac{\pi}{2}$ in all projected subspaces and hence stable in the entire 5D space. We have also checked that for the other class of potentials, the behaviour of the fixed point remains the same. These fixed points represent an accelerating universe with $w_{\phi}=1$ and $w_{tot}=-1$ for $\theta=n\pi$ whereas $w_{\phi}=w_{tot}=-1$ for $\theta=n\pi\pm\frac{\pi}{2}$. Interestingly, in the first case, the scalar field behaves as a stiff fluid.
Figure 2: Phase plots in different $3D$ projected spaces for the equilibrium point ‘$c$’ and for the potential $V_{2}$. The location of the fixed point ‘$c$’ is depicted by the blue point/line. The points/line in magenta color represents the location of the other fixed points which are unstable in nature.
Figure 3: Plots of the numerical solutions for each dynamical variable by considering perturbations around the equilibrium point ‘$ec1$’. The plots in the left column are shown for potential $V_{1}$ and on the right it is shown for the potential $V_{8}$.
@@ V.5 The Era Shared by Scalar Field and other components
The fixed points ‘$er1_{\pm}$’ to ‘$er6_{\pm}$’ represent the state where the universe’s energy budget is shared by radiation and the scalar field by the relation $r_{c}^{2}+\Omega_{rc}=1$. For each $n\in\mathbb{Z}$, the equilibrium points ‘$er1_{\pm}$’, ‘$er2_{\pm}$’, ‘$er5_{\pm}$’, and ‘$er6_{\pm}$’ exhibit positive eigenvalues, making these points unstable or saddle depending on the choice of parameters $\alpha_{0},\alpha_{1},\alpha_{2}$. The eigenvalues of the Jacobian matrix at ‘$er3_{\pm}$’ and ‘$er4_{\pm}$’ are complicated in general, making it difficult to draw any conclusions about their stability. For these points, we evaluate the stability at specific values of $\alpha$ parameters. We have considered the same value of the $\alpha$ parameters that have been used to plot Fig.4 and Fig.6, for potentials $V_{1}$ and $V_{3}$ respectively. For the potential $V_{1}$ the eigenvalues at ‘$er3_{\pm}$’ are $1,4,-2.8394$, and $-0.5\pm 1.93041i$, while at ‘$er4_{\pm}$’ they are $-201.793,-11.2262,10.2262,4$, and $1$, indicating that these points are saddle points. For potential $V_{3}$, the eigenvalues at ‘$er3_{\pm}$’ and ‘$er4_{\pm}$’ are $1,4,-4.48$, and $-0.5\pm 1.92544i$, showing that they are saddle points. The fixed points ‘$er1_{\pm}$’ and ‘$er2_{\pm}$’ exist for the potentials $V_{1}$ and $V_{3}$; ‘$er5_{\pm}$’ exist for the potential $V_{2}$; and ‘$er6_{\pm}$’ for potential $V_{4}$. For all possible equilibrium points in this era $w_{\phi}=w_{tot}=1/3$ representing a decelerated universe.
@@
For the equilibrium points in the era shared by scalar field and matter, we have $r_{c}^{2}+\Omega_{mc}=1$. In this case, the equilibrium points are given by ‘$em1_{\pm}$’ - ‘$em7$’. For each $n\in\mathbb{Z}$, though the eigenvalues of the Jacobian matrix evaluated at the equilibrium points ‘$em1_{\pm}$’, ‘$em2_{\pm}$’ and ‘$em4$’ - ‘$em7$’ depend on the parameters $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$; but it has a positive eigenvalue irrespective of $\alpha_{i}$. So, the equilibrium points given in this era are unstable in nature. The equilibrium points ‘$em1_{\pm}$’ - ‘$em3_{\pm}$’ exist for the potentials $V_{1}$, $V_{3}$; ‘$em4$’ and ‘$em5$’ exist for the potential $V_{2}$; ‘$em6$’ and ‘$em7$’ exist for the potential $V_{4}$. The eigenvalues of the Jacobian matrix evaluated at the equilibrium point ‘$em3_{\pm}$’ are very complicated. So, for ‘$em3_{\pm}$’, we evaluate the stability similarly as the case for ‘$er3_{\pm}$’ and ‘$er4_{\pm}$’. For the potential $V_{1}$ the eigenvalues at ‘$em3_{+}$’ are $3,-1,-2.12955$, and $-0.75\pm 1.97931i$, while at ‘$er3_{-}$’ they are $9.06334,3,-1,-10.5633$, and $-151.345$, indicating that these points are saddle points. For potential $V_{3}$, the eigenvalues at ‘$em3_{\pm}$’ are $3,-1,-3.36$, and $-0.75\pm 1.97522i$, showing that they are saddle points. The exact value of the deceleration parameter depends on $\alpha$ parameters for all the potentials belonging to this era. However, for all possible equilibrium points in this era $w_{\phi}=w_{tot}=0$ or $1/3$ representing a decelerated universe.
@@
For the equilibrium points in the era shared by scalar field and $\Lambda$, we have $r_{c}^{2}+\Omega_{\Lambda_{c}}=\Omega_{\phi_{c}}+\Omega_{\Lambda_{c}}=1$. In this case, the equilibrium points are given by ‘$ec1$’ and ‘$ec2_{\pm}$’. The equilibrium points given by ‘$ec1$’ exist for all potentials $V_{1}-V_{8}$. The nature of the equilibrium points given by ‘$ec1$’ depends on the parameter $\alpha_{0}$. If $0<\alpha_{0}<\frac{3}{4r_{c}^{2}}$, then the equilibrium points given by ‘$ec1$’ are normally hyperbolic equilibrium points and the Jacobian matrix has four eigenvalues with negative real parts and one zero eigenvalue. Again, if $\alpha_{0}=0$ then the equilibrium points given by ‘$ec1$’ are of non-hyperbolic type and the Jacobian matrix has three negative and two zero eigenvalues. To check the stability of those equilibrium points we plot the evolution of the dynamical variables $r,\theta,\Omega_{m},\Omega_{r}$ and $\lambda$ by solving the system numerically where the system has been perturbed from the equilibrium point ‘$ec1$’. This approach for finding the stability of the non-hyperbolic fixed points has been previously used in [78, 80, 81]. In Fig.3 for example we have plotted the evolution of the system with $N$ for the potentials $V_{1}$ (on the left column) and $V_{8}$ (on the right column) subjected to perturbations around the fixed point ‘$ec_{1}$’. We can see that the evolution of the dynamical variables does not diverge but rather stays in the neighbourhood of the fixed point, indicating its stable nature. We have also checked that the qualitative behaviour of the fixed point remains the same even for the other potentials. Thus, the fixed points given by ‘$ec1$’ are stable for $\alpha_{0}=0$. Again, if $\alpha_{0}<0$, then the equilibrium points given by ‘$ec1$’ are non-hyperbolic type and the Jacobian matrix has a positive eigenvalue. Thus, the equilibrium points given by ‘$ec1$’ are stable for $0\leq\alpha_{0}\leq\frac{3}{4r_{c}^{2}}$, where $0\leq r_{c}\leq 1$ and unstable for $\alpha_{0}<0$. These fixed points represent an accelerating universe with $w_{\phi}=w_{tot}=-1$.